# Primitive-Root Determinant Densities over Prime Fields and Implications for PRIM-LWE

### 21 Apr 2026 v4

@@
Abstract
For a prime $p$, let $c_{n}(p)$ denote the fraction of $n\times n$ matrices over $\mathbb{F}_{p}$ whose determinant is a primitive root modulo $p$, and let $c(p)=\lim_{n\to\infty}c_{n}(p)$ be the limiting density. This quantity governs the efficiency of PRIM-LWE, a variant of the learning with errors (LWE) problem in which the secret matrix is required to have primitive-root determinant. The value $1/c(p)$ equals the expected rejection-sampling overhead in the reduction from standard LWE to PRIM-LWE. We prove unconditionally that $\inf_{p\text{ prime}}c(p)=0$, resolving an open question from the PRIM-LWE literature, and establish the rate $\min_{p\leq x}c(p)\asymp 1/\log\log x$ as $x\to\infty$. We show that $c(p)$, viewed as a function on the primes, has a continuous, purely singular limiting distribution with support $[0,1/2]$ and Hausdorff dimension 0. The moments of this limiting distribution are given by convergent Euler products, and the Mellin transform $\mathbb{E}[X^{s}]$ extends analytically to $\operatorname{Re}(s)>0$. Using the number of distinct prime factors $\omega(p-1)$, we derive explicit lower bounds: for every prime $p>2^{30}$ the overhead satisfies $1/c(p)<1.79\log p$, and in general $1/c(p)=O(\log\log p)$. We also describe a certified algorithm for evaluating $c(p)$ and tabulate values for primes of cryptographic interest.
Keywords: Primitive root, determinant density, PRIM-LWE, limiting distribution, Hausdorff dimension
MSC 2020: 11A07, 11N60, 94A60, 15B33
1 Introduction
Let $p$ be a prime and $M_{n}(\mathbb{F}_{p})$ denote the ring of $n\times n$ matrices over $\mathbb{F}_{p}$. Following [36], a matrix $M\in M_{n}(\mathbb{F}_{p})$ has primitive-root determinant if $\det(M)$ generates $\mathbb{F}_{p}^{*}$. Write $c_{n}(p)$ for the fraction of such matrices, and $c(p):=\inf_{n\geq 1}c_{n}(p)=\lim_{n\to\infty}c_{n}(p)$ for the limiting density; the precise formulas are recorded in Section 2. Since $c_{n}(p)$ is decreasing in $n$, one has $c(p)\leq c_{n}(p)$ for every $n\geq 1$. Consequently, every lower bound on $c(p)$ is simultaneously a lower bound on every finite-dimensional density $c_{n}(p)$.
The quantity $c(p)$ arises as a reduction constant in the PRIM-LWE variant [36] of the popular learning with errors (LWE) problem [33], introduced to construct verifiable secret sharing schemes with “free” verification. The reduction proceeds by rejection-sampling and succeeds with probability $c_{n}(p)$ in dimension $n$, so the dimension-uniform overhead is $1/c(p)$. This overhead represents a tightness loss in the security reduction: the hardness guarantee for PRIM-LWE degrades by a factor of $c(p)$ relative to standard LWE, making tight bounds on $c(p)$ essential for concrete security estimates. Sehrawat, Yeo, and Desmedt [36, Remark 3(ii)] asked whether $\inf_{p}c(p)=0$ can be proved unconditionally.
@@ 1 Introduction
1. Unconditional vanishing (Theorem 2). We prove $\inf_{p\text{ prime}}c(p)=0$ using only Dirichlet’s theorem on primes in arithmetic progressions and Mertens’ third theorem, without assuming the primorial-prime hypothesis; see Section 3.
2. Sharp order of the minimum (Proposition 1, Corollary 7). We establish
@@
See Remark 4 for effectivity.
3. Limiting distribution (Theorems 4 and 5, Corollary 2, Proposition 4). The quantity $c(p)$ possesses a continuous, purely singular limiting distribution over the primes whose support is exactly $0,\tfrac{1}{2}]$; the limiting measure has Hausdorff dimension 0. The full-support conclusion is new; it combines the left non-differentiability result of Deshouillers–Hassani [9, Theorem 1.1] with the density of even totient ratios (Lemma 4). The moments of $G$ are given by convergent Euler products (Theorem 6), the Mellin transform $\mathbb{E}[X^{s}]$ extends analytically to $\{\operatorname{Re}(s)>0\}$ (Theorem 8), and $G$ is not Hölder of any positive order at $1/2$ (Proposition 7).
Since the limiting law is purely singular and supported on a set of Hausdorff dimension 0, it cannot certify $c(p)$ at any individual prime; this motivates the pointwise bounds of item 4.
4. Explicit bounds via distinct prime factors (Theorem 9). We give unconditional lower bounds on $c(p)$ for primes $p$ parameterized by $\omega(p-1)$, the number of distinct prime factors of p1p-1, yielding computable bounds for any prime $p$ such that the factorization of p1p-1 is known. For instance, for the NIST-standardized primes $p=3329$ and $p=8380417$, the exact factorizations of p1p-1 give $c(3329)\geq 0.4613$ and $c(8380417)\geq 0.2932$; see Section 7. In particular, the dimension-uniform expected rejection-sampling overhead $1/c(p)$ in the LWE-to-PRIM-LWE reduction [36, Theorem 10] admits an explicit upper bound depending only on $\omega(p-1)$.
From the computational viewpoint, our main contributions are: explicit lower bounds on $c(p)$ valid for all primes $p>2^{30}$ (Corollary 9), a certified rational-arithmetic algorithm for evaluating $c(p)$ to arbitrary precision given the factorization of p1p-1 (Proposition 8), and tables of $c(p)$ for all primes appearing in current NIST post-quantum cryptographic standards and representative lattice-based implementations (Tables 2 and 3). The explicit bound takes the following compact form:
Corollary 1 (Cryptographic consequence).
Let $P_{3}:=\prod_{j=1}^{\infty}(1-1/3^{j})\approx 0.5601260779$. If $p\geq 3$ is a prime modulus with $\omega(p-1)\leq K$, then
$\frac{1}{c(p)}\;\leq\;\frac{1}{P_{3}}\prod_{i=1}^{K}\frac{p_{i}}{p_{i}-1}.$
The general $\omega(p-1)$-based bound is immediate from Theorem 9. Table 1 lists representative values of $B(K)$, and Tables 2 and 3 collect these bounds for representative cryptographic moduli.
Related work
A related distribution problem was studied by Aivazidis and Sofos [1], who considered the density of maximal-order elements in $\mathrm{GL}_{n}(q)$ as $q$ varies over prime powers and proved corresponding distribution laws. Their setting is different from ours: we work over prime fields and with the primitive-root determinant density arising in PRIM-LWE. On the distributional side, Deshouillers and Hassani [9, Proposition 2.1 and equation (2.1)] proved limiting-distribution laws for $\varphi(p-1)/(p-1)$ in the residue classes $p\equiv 1\pmod{m}$, including continuity, and [9, Theorem 1.1] established left non-differentiability of the distribution function at every value $\varphi(m)/m$ with $m$ even. The analogous support question for the classical distribution of $\varphi(n)/n$ over the integers remains open; see [9].
2 Background and the Open Question
@@
Throughout, $\log$ denotes the natural logarithm. We write $\varphi$ for Euler’s totient function, $\omega(n)$ for the number of distinct prime factors of $n$, and $p_{k}$ for the $k$-th prime.
Following [36], define
@@
(see, e.g., Mullen and Panario [30, Theorem 13.3.3]), it follows that
@@
cf. [36, Corollary 4]. Consequently, the limit
@@
Writing $p-1=r_{1}^{e_{1}}\cdots r_{k}^{e_{k}}$ for the prime factorization of p1p-1, one has
$c(p)\;=\;\prod_{i=1}^{k}\Bigl(1-\frac{1}{r_{i}}\Bigr)\;\cdot\;\prod_{j=1}^{\infty}\Bigl(1-\frac{1}{p^{j}}\Bigr).$ (1)
Since $P(p):=\prod_{j=1}^{\infty}(1-p^{-j})\to 1$ as $p\to\infty$, the behavior of $c(p)$ over the primes is governed by $\varphi(p-1)/(p-1)$, and the distributional theory in Section 5 reduces to the classical shifted-prime distribution of this ratio.
In [36, Remark 3(ii)], Sehrawat, Yeo, and Desmedt observed that if there exist infinitely many primorial primes $p=\prod_{i=1}^{k}p_{i}+1$, then $\varphi(p-1)/(p-1)=\prod_{i=1}^{k}(p_{i}-1)/p_{i}\to 0$ by Mertens’ third theorem, which would imply $\inf_{p}c(p)=0$. They left open whether this can be proved unconditionally; see [7] for computational results on primorial primes.
3 Unconditional Proof that $\inf_{p}c(p)=0$
@@
Theorem 1 (Mertens’ Third Theorem [28]; see [29, Theorem 2.7(e)], 1874).
@@
Lemma 1 (Lower bound for the matrix-count Euler product).
@@
Theorem 2 (Unconditional vanishing of $\inf c(p)$).
@@
By Mertens’ third theorem (Theorem 1),
@@ 4 Quantitative Refinements
Lemma 2 (Prime-number-theorem consequences for primorials; cf. [29, Corollary 2.5]).
@@
Proposition 1 (Upper bound for the minimum density up to $x$).
@@
Since $\log N_{k}=\sum_{i=1}^{k}\log p_{i}\sim p_{k}$ by the prime number theorem, the inequality $N_{k}^{\,L}\leq x$ gives
$L\,\log N_{k}\leq\log x,$
@@
By construction the least prime $p^{*}(k)\equiv 1\pmod{N_{k}}$ satisfies $p^{*}(k)\leq x$, and the argument in the proof of Theorem 2 gives
@@
The matching lower bound $\min_{p\leq x,\,p\text{ prime}}c(p)\gg 1/\!\log\log x$, which completes the proof that the order is sharp, is established in Corollary 7 below; its proof requires only the primorial counting argument and Lemma 1.
5 The Limiting Distribution of $c(p)$
The classical existence and continuity of the limiting distribution of $\varphi(p-1)/(p-1)$ over primes follow from the shifted-prime Erdős–Wintner framework [16] (Tenenbaum [38, pp. 262–263], Deshouillers–Hassani [9]). Throughout this section we write
$f(n)\;:=\;\log\frac{n}{\varphi(n)}\;=\;\sum_{\begin{subarray}{c}\ell\mid n\\ \ell\textup{ prime}\end{subarray}}\log\frac{\ell}{\ell-1}$ (2)
5.1 The distribution of $\varphi(p-1)/(p-1)$
Lemma 3 (Verification of the Erdős–Wintner hypotheses).
The function $f$ defined in (2) is strongly additive, with $f(n)=\log(n/\varphi(n))=-\log(\varphi(n)/n)$. For every prime $\ell$,

### 06 May 2026 v5

# Primitive-root ratio over prime fields: a shifted-prime distribution of Hausdorff dimension zero and Implications for PRIM-LWE

@@
Abstract.
For a prime $p$, let $c(p)$ denote the limiting fraction of $n\times n$ matrices over $\mathbb{F}_{p}$ whose determinant is a primitive root modulo $p$. The quantity $c(p)$ is a natural multiplicative deformation of the totient ratio $\varphi(p-1)/(p-1)$ and inherits its distributional behaviour over the primes. Existence and continuity of the limiting law follow from the shifted-prime Erdős–Wintner–Hildebrand framework. We prove the following new results: unconditionally, $\inf_{p}c(p)=0$ and the sharp order is $\min_{p\leq x}c(p)\asymp 1/\log\log x$; the reciprocal satisfies $\limsup_{p\to\infty,\,p\textup{ prime}}1/(c(p)\log\log p)=e^{\gamma}$, and no smaller constant suffices. We give a complete proof, combining an adaptation of Erdős’s argument with the Jessen–Wintner pure-type dichotomy, that the limiting distribution is purely singular, and strengthen this to $\dim_{H}(\mu_{G})=0$, i.e. the limiting measure is carried by a Borel set of Hausdorff dimension zero. The distribution has full topological support $0,\tfrac{1}{2}]$ and admits a Bernoulli-product representation indexed by the odd primes. The moments are given by convergent Euler products, and the Mellin transform $\mathbb{E}[X^{s}]$ extends to an entire function of $s$, non-vanishing on $\operatorname{Re}(s)>0$. Near the right endpoint, $1-G(\tfrac{1}{2}-\varepsilon)\sim\kappa/\log(1/\varepsilon)$ with an explicit constant $\kappa$. The quantity $1/c(p)$ equals the dimension-uniform expected rejection-sampling overhead in the reduction from learning with errors (LWE) to PRIM-LWE in lattice-based cryptography. The explicit bounds yield concrete overhead estimates for all primes appearing in current NIST post-quantum standards and representative NTT-friendly moduli.
2020 Mathematics Subject Classification:
11N60, 11A07, 11K65, 28A78, 94A60
1. Introduction
The distribution of the Euler totient ratio $\varphi(n)/n$ over the positive integers was first investigated by Schoenberg [30], who showed that the sequence possesses a continuous asymptotic distribution function. The existence of such a distribution was placed on a rigorous probabilistic footing by the Erdős–Wintner theorem [11], which identifies the class of real additive functions admitting a limiting law. Erdős [10] subsequently proved that this distribution is purely singular with respect to Lebesgue measure, and the Jessen–Wintner zero–one law [18] supplies the underlying dichotomy. The shifted-prime analogue—the behaviour of $\varphi(p-1)/(p-1)$ as $p$ ranges over primes—was addressed by Kátai [20] and, in definitive form, by Hildebrand [16], who extended the Erdős–Wintner framework to additive functions evaluated on shifted primes. Deshouillers and Hassani [7] refined this picture by establishing continuity in arithmetic progressions and proving left non-differentiability of the distribution function at every value $\varphi(m)/m$ with $m$ even.
The quantity
$c(p)\;:=\;\frac{\varphi(p-1)}{p-1}\;\prod_{j=1}^{\infty}\Bigl(1-\frac{1}{p^{j}}\Bigr)$
arises as the limiting density of $n\times n$ matrices over $\mathbb{F}_{p}$ whose determinant is a primitive root modulo $p$; the precise derivation is recalled in Section 2. Writing $P(p):=\prod_{j=1}^{\infty}(1-p^{-j})$, one has $P(p)\to 1$ as $p\to\infty$, so the behaviour of $c(p)$ is governed by the classical totient ratio $\varphi(p-1)/(p-1)$, and the distributional theory of $c(p)$ reduces to the shifted-prime Erdős–Wintner–Hildebrand framework. In this sense $c(p)$ is a natural multiplicative deformation of $\varphi(p-1)/(p-1)$.
The open question whether $\inf_{p}c(p)=0$ was raised in [31, Remark 3(ii)]; the cryptographic implications of the explicit bounds are developed in Section 8.
@@ 1. Introduction
We consolidate our principal findings into four theorems.
[…]
@@
The extremal order of the reciprocal is sharp:
$\limsup_{p\to\infty,\,p\textup{ prime}}1/(c(p)\,\log\log p)=e^{\gamma}\approx 1.781.$
In particular, $1/c(p)\leq(e^{\gamma}+o(1))\log\log p$ for every prime $p$, and no smaller constant suffices. The limsup is realised asymptotically along the least primes $p_{k}^{*}\equiv 1\pmod{N_{k}}$, whose existence is guaranteed by Dirichlet’s theorem and whose size is controlled by Linnik’s theorem.
See Theorem 3.3, Proposition 4.2, Corollary 7.2, and Theorem 7.4. Remark 7.5 discusses effectivity.
Theorem B (Limiting law).The quantity $c(p)$ possesses a continuous limiting distribution $G$ over the primes, coinciding with the distribution of $\varphi(p-1)/(p-1)$. The topological support of $G$ is the full interval $0,\tfrac{1}{2}]$, and $G$ admits a Bernoulli-convolution representation: if $B_{\ell}$ are independent Bernoulli random variables with $\mathbb{P}(B_{\ell}=1)=1/(\ell-1)$ for each odd prime $\ell$, then
$X\;\stackrel{{\scriptstyle d}}{{=}}\;\tfrac{1}{2}\prod_{\ell\geq 3,\;\ell\textup{ prime}}\Bigl(1-\frac{1}{\ell}\Bigr)^{\!B_{\ell}}.$
See Theorems 5.5 and 5.8, Corollary 5.14, and Lemma 6.1. Equivalently, the logarithmic transform of the distribution is an infinite convolution of two-point measures (Lemma 6.1).
Theorem C (Hausdorff dimension zero and singularity).The Hausdorff dimension of the limiting measure $\mu_{G}$ is zero: $\dim_{H}(\mu_{G})=0$. In particular, $\mu_{G}$ is purely singular with respect to Lebesgue measure.
See Theorem 6.2 and Section 5.3.
Theorem D (Mellin transform and endpoint law).For every $s\in\mathbb{C}$, the Mellin transform $M(s)=\mathbb{E}[X^{s}]$ is given by a convergent Euler product; it extends to an entire function of $s$, non-vanishing on $\{\operatorname{Re}(s)>0\}$. Near the right endpoint, $1-G(\tfrac{1}{2}-\varepsilon)\sim\kappa/\log(1/\varepsilon)$ as $\varepsilon\downarrow 0$, where $\kappa=\mathfrak{S}_{2}\,e^{-\gamma}$ and $\mathfrak{S}_{2}$ is the Hardy–Littlewood twin-prime singular series.
See Theorems 6.3 and 6.4, Proposition 6.9, and Corollary 6.10. The twin-prime singular series enters because the event that no small prime $\ell$ divides p1p-1 forces $\varphi(p-1)/(p-1)$ close to $1/2$. The resulting local factors $1-1/(\ell-1)^{2}$ are precisely those governing the density of twin primes in the Hardy–Littlewood heuristic [15].
As a consequence, explicit lower bounds on $c(p)$ via $\omega(p-1)$ yield effective estimates for any prime whose factorisation of p1p-1 is known; see Section 7.
Davenport [6] independently studied the distribution of $\sigma(n)/n$ over the integers, where $\sigma$ denotes the sum-of-divisors function; for $\varphi(n)/n$ the topological support is the full interval $[0,1]$ and the global modulus of continuity remains unknown (see Section 9). Hildebrand’s proof [16] of the shifted-prime Erdős–Wintner theorem builds on the equidistribution of primes in arithmetic progressions on average established by Bombieri [3]; see also Tenenbaum [32, 33]. Tenenbaum [32] remarked that the pure singularity of the shifted-prime distribution can be obtained by adapting an exercise solution in [34, Exercise 2.4] (corresponding to [33, Exercise 256]), and Deshouillers–Hassani [7] cite the same exercise, but neither source carries out the argument. We give a complete proof in Section 5.3 below, combining an explicit adaptation of Erdős’s argument with the Jessen–Wintner pure-type dichotomy [18]. We strengthen this to the metric result $\dim_{H}(\mu_{G})=0$ in Section 6.1. Wintner’s convergence theory [36]—the founding document for infinite convolutions of discrete measures—supplies the underlying framework. The Hausdorff dimension of Bernoulli convolutions has a substantial literature; see Kahane [19] for an early observation that Erdős-type arguments yield Hausdorff dimension estimates, and Peres, Schlag, and Solomyak [26] and Peres–Solomyak [27] for further dimensional results. The classical Bernoulli convolution involves a fixed contraction ratio. The infinite convolution arising here involves varying ratios $\log(\ell/(\ell-1))$ indexed by the odd primes, so the standard Pisot–Salem machinery does not apply directly. Billingsley [2] and Elliott [13] provide modern probabilistic perspectives.
The extremal behaviour of $\varphi(n)/n$ dates to Mertens [23], whose third theorem gives $\prod_{\ell\leq x}(1-1/\ell)\sim e^{-\gamma}/\log x$; the systematic study of extremal values of arithmetic functions was initiated by Ramanujan [28]. Over the primes, primes $p\equiv 1\pmod{N_{k}}$—for which p1p-1 has many small prime factors—realise the smallest values of the totient ratio, and the size of the least such prime is controlled by Linnik’s theorem [21]. Aivazidis and Sofos [1] considered the density of maximal-order elements in $\mathrm{GL}_{n}(q)$ over prime powers. There is also a conceptual kinship with the Artin primitive-root conjecture (see Hooley [17] for a conditional proof under GRH): while Artin’s conjecture concerns the behaviour of a fixed base as $p$ varies, the present work concerns the aggregate density $c(p)$ in the complementary regime where the base varies uniformly. Both problems ultimately depend on the factorisation of p1p-1.
Organisation
Section 2 recalls the derivation of $c(p)$ and the open question from [31]. Section 3 resolves this question unconditionally, and Section 4 gives quantitative refinements via Linnik’s theorem. Section 5 develops the distributional theory: existence and continuity of the limiting law (Section 5.1), transfer from $\varphi(p-1)/(p-1)$ to $c(p)$ (Section 5.2), and the complete proof of pure singularity (Section 5.3). Section 6 establishes the convolution structure and Hausdorff dimension zero (Section 6.1), the moment Euler product and Mellin transform (Section 6.2), the Fourier structure (Section 6.3), and the endpoint asymptotic (Section 6.4). Section 7 provides explicit lower bounds, Section 8 applies these bounds to the PRIM-LWE reduction in lattice-based cryptography, and Section 9 concludes by discussing two open questions.
2. Background and the Open Question
@@
Throughout, $\log$ denotes the natural logarithm. We write $\omega(n)$ for the number of distinct prime factors of $n$ and $p_{k}$ for the $k$-th prime.
Define (cf. [31])
@@
(see, e.g., Mullen and Panario [25, Theorem 13.3.3]), it follows that
@@
cf. [31, Corollary 4]. Consequently, the limit
@@
Writing $p-1=r_{1}^{e_{1}}\cdots r_{k}^{e_{k}}$ for the prime factorisation of p1p-1, one has
(2.1) $c(p)\;=\;\prod_{i=1}^{k}\Bigl(1-\frac{1}{r_{i}}\Bigr)\;\cdot\;\prod_{j=1}^{\infty}\Bigl(1-\frac{1}{p^{j}}\Bigr).$
Since $P(p):=\prod_{j=1}^{\infty}(1-p^{-j})\to 1$ as $p\to\infty$, the behaviour of $c(p)$ over the primes is governed by $\varphi(p-1)/(p-1)$, and the distributional theory in Section 5 reduces to the classical shifted-prime distribution of this ratio.
In [31, Remark 3(ii)], Sehrawat, Yeo, and Desmedt observed that if there exist infinitely many primorial primes $p=\prod_{i=1}^{k}p_{i}+1$, then $\varphi(p-1)/(p-1)=\prod_{i=1}^{k}(p_{i}-1)/p_{i}\to 0$ by Mertens’ third theorem, which would imply $\inf_{p}c(p)=0$. They left open whether this can be proved unconditionally; see [4] for computational results on primorial primes.
3. Unconditional Proof that $\inf_{p}c(p)=0$
@@
Theorem 3.1 (Mertens’ Third Theorem [23]; see [24, Theorem 2.7(e)], 1874).
@@
Lemma 3.2 (Lower bound for the matrix-count Euler product).
@@
Theorem 3.3 (Unconditional vanishing of $\inf c(p)$).
@@
By Mertens’ third theorem (Theorem 3.1),
@@ 4. Quantitative Refinements
Lemma 4.1 (Prime-number-theorem consequences for primorials; cf. [24, Theorem 6.9 and Corollary 2.5]).
@@
Proposition 4.2 (Upper bound for the minimum density up to $x$).
@@
Since $\log N_{k}=\sum_{i=1}^{k}\log p_{i}\sim p_{k}$ by the prime number theorem, the inequality $C_{0}N_{k}^{\,L}\leq x$ gives
$L\,\log N_{k}\leq\log x-\log C_{0},$
@@
By construction the least prime $p^{*}(k)\equiv 1\pmod{N_{k}}$ satisfies $p^{*}(k)\leq x$, and the argument in the proof of Theorem 3.3 gives
@@
The matching lower bound $\min_{p\leq x,\,p\text{ prime}}c(p)\gg 1/\!\log\log x$, which completes the proof that the order is sharp, is established in Corollary 7.2; its proof requires only the primorial counting argument and Lemma 3.2. In particular, the lower bound is independent of the distributional machinery in Sections 56: it uses the explicit inequality $c(p)\geq P_{3}\prod_{i=1}^{\omega(p-1)}(1-1/p_{i})$ of Theorem 7.1 together with the primorial comparison $\omega(p-1)\leq m(p-1)$ and Mertens’ third theorem.
5. The Limiting Distribution of $c(p)$
The classical existence and continuity of the limiting distribution of $\varphi(p-1)/(p-1)$ over primes follow from the shifted-prime Erdős–Wintner framework [11] (Tenenbaum [32], Deshouillers–Hassani [7]; see also [34]). We recall these results in Sections 5.1 and 5.2 for completeness. Pure singularity of the shifted-prime law was anticipated by a remark of Tenenbaum [32], who noted that the integer-case argument of Erdős [10] (with solution in [34, Exercise 2.4]) can be adapted. We carry out this adaptation in Section 5.3, providing a complete proof in the shifted-prime setting. The strictly stronger result $\dim_{H}(\mu_{G})=0$ is established in Section 6.1. Throughout this section we write
(5.1) $f(n)\;:=\;\log\frac{n}{\varphi(n)}\;=\;\sum_{\begin{subarray}{c}\ell\mid n\\ \ell\textup{ prime}\end{subarray}}\log\frac{\ell}{\ell-1}\,.$
5.1. The distribution of $\varphi(p-1)/(p-1)$
Lemma 5.1 (Verification of the Erdős–Wintner hypotheses).
The function $f$ defined in (5.1) is strongly additive, with $f(n)=\log(n/\varphi(n))=-\log(\varphi(n)/n)$. For every prime $\ell$,